# NOBLE - Neural Operator with Biologically-informed Latent Embeddings to Capture Experimental Variability in Biological Neuron Models

### v2

@@ 3.2 Current Amplitude Sampling
We sample square-pulse amplitudes from a skew-normal distribution whose support matches the experimentally validated range of our HoF models, $I\in[-0.11,0.28]\mathrm{nA}$. To effectively capture the highly nonlinear dynamics around the spiking threshold (00 to $0.05\mathrm{nA}$) where neural responses transition abruptly from being non-spiking to spiking, the mode of our sampling distribution is strategically located within this peri-threshold window. To address the greater learning challenge posed by the high-frequency components of depolarizing, spiking responses (characterised by features such as spike width, latency to first spike, and spike count), we deliberately use a heavier positive tail in our sampling distribution. This bias ensures the model encounters numerous examples of spike onset and complex spiking patterns during training while still covering the full validated input range. The distribution of square-pulse amplitudes considered is shown in the appendix in Figure 11.
@@ 3.3 Embedding Strategy for Neuron-Model Variability
The proposed NOBLE framework learns a single neural operator that maps from a continuous latent space of user-defined, interpretable neuron characteristics to an ensemble of somatic voltage responses induced by current injection. For the PVALB neuron considered in our numerical experiments, the firing rate (F-I) curve is a useful electrophysiological descriptor that summarises cellular excitability by relating injected current amplitude to the neuron’s firing rate [47]. Differences between HoF parameterisations manifest as shifts in key features of this curve: the threshold current $I_{\textit{thr}}$ (the minimum amplitude that elicits spiking) and the local slope $s_{\textit{thr}}$ at $I_{\textit{thr}}$, which measures the rate of increase in the firing rate upon spiking. Figure 5A displays examples of F-I curves for different HoF PVALB models, showing the variability in $I_{\textit{thr}}$ and $s_{\textit{thr}}$ across models.
Using this observation, we propose representing a given PVALB neuron model by its threshold current $I_{\textit{thr}}$ and local slope $s_{\textit{thr}}$, that is, we represent the model as $(I_{\textit{thr}},s_{\textit{thr}})$. We propose to use this representation as part of a NeRF-style (Neural Radiance Field) embedding [48], where input features are encoded using sine and cosine functions. More precisely, a feature $p$ is encoded as a stack of trigonometric time-series
$\left[\sin(2^{0}\pi p\odot t),\ \cos(2^{0}\pi p\odot t),\ \ldots,\ \sin(2^{K-1}\pi p\odot t),\ \cos(2^{K-1}\pi p\odot t)\right]$ (2)
@@
In our PVALB neuron example, we encode separately the model features $I_{\textit{thr}}$ and $s_{\textit{thr}}$, and the current injection, using this strategy and stack the resulting embeddings along the channel dimension of the input to the FNO. To ensure periodicity and promote distant representations of distinct HoF models in the feature space, we normalize both $I_{\textit{thr}}$ and $s_{\textit{thr}}$ to the range $[0.5,3.5]^{2}$ before embedding. Figure 3 displays the latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of the 60 HoF PVALB models used in our numerical experiments. This figure also shows that using only one of $I_{\textit{thr}}$ and $s_{\textit{thr}}$ instead of both would not allow to distinguish very well between different HoF models.
@@ 3.4 NOBLE: Neural Operator with Biologically-informed Latent Embeddings
We introduce the Neural Operator with Biologically-informed Latent Embeddings (NOBLE), for modeling neuronal voltage dynamics in response to current injections. NOBLE offers a scalable alternative to computationally intensive numerical solvers for biophysically detailed, PDE-based neuron models. It learns a direct mapping from input currents and an interpretable, continuous latent space of neuron features to voltage traces (Figure 1A). A key feature of NOBLE is its use of biologically-informed embeddings to generalize across variability in biological neuron models.
At its core, NOBLE is based on a temporal neural operator. The discretization invariance of neural operators allows for NOBLE to learn on low-resolution data and infer somatic voltage dynamics at higher resolutions. By combining a neural operator with the proposed continuous interpretable embedding, NOBLE learns a continuous operator over the space of bio-realistic neuron models.
The proposed NOBLE framework offers several key advantages that set it apart from existing approaches:
  • While prior works model neuronal variability using many models, each generating a single prediction, NOBLE proposes a unifying framework combining all such models into one. Instead of training separate models for each experimentally observed neuronal variation, NOBLE learns directly the ensemble dynamics with the neuronal variability. Conditioned on a particular electrophysiological feature, NOBLE generates one realization of the variability in a biological neuron model.
  • NOBLE can interpolate between model features in its embedding space to generate new bio-realistic neuron models. This is a striking property that no prior work offers, (e.g. interpolation between parameters of PDE models offers no guarantee of validity for the resulting model).
  • With the proposed continuous embedding space, NOBLE can be used to generate infinitely many different instances of neuron dynamics (both spiking and subthreshold), since it can interpolate between model features to generate new neuron models.
  • The biologically-informed latent space of NOBLE, combined with its ability to generate arbitrarily many biophysically realistic neuron models, enables fine-grained exploration of neural dynamics. By sampling models across this space, NOBLE can produce somatic responses to diverse inputs, allowing for visualization techniques, such as heat maps and surface plots, to reveal how latent features influence electrophysiological behavior.
We demonstrate these novel characteristics of NOBLE with extensive numerical experiments, next.
@@ 4.1 Experimental Setup
Figure 3: Latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of $\{\text{HoF}_{i}^{\textit{train}}\}$ (black dots) and $\{\text{HoF}_{i}^{\textit{test}}\}$ (red crosses) models. The latter lie in the convex hull $\mathcal{CH}_{\textit{train}}$ of $\{\text{HoF}_{i}^{\textit{train}}\}$.
@@
We consider the PVALB neuron modeling example introduced in Section 2.1. Given a PVALB model $\text{HoF}_{\ell}$ and a current injection $I$, the neural operator in NOBLE takes the stacked embeddings of the (normalized) model features $I_{\textit{thr}}$ and $s_{\textit{thr}}$ associated to $\text{HoF}_{\ell}$ (with $K=1$ frequency) and $I$ (with $K=9$ different frequencies), and outputs the somatic voltage response. We have access to 60 HoF models $\{\text{HoF}_{i}\}_{i=1}^{60}$ obtained using a multi-objective evolutionary optimization strategy. We use 50 of them during training, $\{\text{HoF}_{i}^{\textit{train}}\}_{i=1}^{50}$, and keep the remaining 10 HoF models $\{\text{HoF}_{i}^{\textit{test}}\}_{i=1}^{10}$ for testing. Figure 3 displays the latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of these HoF models, and in particular distinguishes between those included and excluded during training. The training dataset is composed of $37,800$ samples, where the current injections are sampled as described in Section 3.2, each of which is associated randomly to one of $\{\text{HoF}_{i}^{\textit{train}}\}$. The samples are generated using a numerical solver [41] built on the NEURON simulation environment [42].
@@
We choose the neural operator in NOBLE to be a 1D FNO with $12$ layers, each with $24$ hidden channels and $256$ Fourier modes, and use the FNO implementation of the NeuralOperator library [46]. The resulting NOBLE with $1.8$M trainable parameters is trained in PyTorch for $350$ epochs to minimize the relative L4 error to the reference HoF simulations, using the Adam optimizer with learning rate $0.008$, and the ReduceLROnPlateau scheduler with factor $0.4$ and patience $4$.
The trained NOBLE generate predictions significantly faster than the reference numerical solver [41]. We record the time necessary to generate $10,000$ predictions on a workstation equipped with a single NVIDIA RTX 4090 GPU (24GB VRAM), an AMD Ryzen 9 7900X CPU, and 64GB of system RAM. The numerical solver only generates a single prediction at a time, and takes roughly $36,\!200$ seconds to generate $10,\!000$ predictions. When doing one inference at a time with NOBLE (i.e. batch size of $1$), we generate $10,\!000$ predictions in $157$ seconds, i.e. a speedup of approximately $230$× compared to the solver. In addition, NOBLE can easily be accelerated on a single GPU by generating multiple predictions at the same time. In particular, with a batch size of $1000$, NOBLE generates $10,\!000$ predictions in $8.59$ seconds, i.e. a speedup of approximately $4200$× compared to the solver.
4.2 Testing on HoF Models Experienced during Training
@@
We begin by validating that the trained NOBLE can accurately replicate the somatic voltage responses of the HoF models $\{\text{HoF}_{i}^{\textit{train}}\}$ used for training, when tested on current injections that have not been experienced during training. As shown in Figure 4A, the voltage traces exhibit minimal differences, confirming that NOBLE works and generalizes well to unseen current injections. This is also confirmed by a L2 relative error of $2.18\%$ on the test set with the $\{\text{HoF}_{i}^{\textit{train}}\}$ models. We further generate the F-I curves using the trained NOBLE for $\{\text{HoF}_{i}^{\textit{train}}\}$, and compare them to the reference F-I curves produced by the numerical solver using the same HoF models. As shown in Figure 5B (top), the curves from both methods remain close, although for 3 out of the 50 HoF models, the NOBLE predictions exhibit higher variations in firing rate for high current amplitude.
Figure 4: Successful predictions of NOBLE for model features experienced during training (In-Distribution) and model features excluded from training (Out-of-Distribution).
@@ 4.3 Interpolating Between Models
Figure 5: F-I curves and distributions of voltage traces for A) experimental versus HoF simulations (PDE) and B) NOBLE predictions versus HoF simulations (PDE). The top row shows the results for $\{\text{HoF}_{i}^{\textit{train}}\}$, while the bottom row shows the results for interpolation between these models.
We now test NOBLE with the neuron model features associated with the HoF models $\{\text{HoF}_{i}^{\textit{test}}\}$ that were not used for training. Note that these models are located within the convex hull $\mathcal{CH}_{\textit{train}}$ of the HoF models $\{\text{HoF}_{i}^{\textit{train}}\}$ used for training, in the embedding space of the PVALB neuron models. Given a model $\text{HoF}_{k}^{\textit{test}}$ excluded during training, we sample randomly 50 points in the model latent space close to the defining $(I_{\textit{thr}},s_{\textit{thr}})$ features of $\text{HoF}_{k}^{\textit{test}}$. This is illustrated in Figure 12 where we show the small neighborhood around a given model in which the 50 samples are sampled. In particular, none of these 50 new models have been experienced during training. Given a current injection, we then run the 50 corresponding NOBLE inferences in parallel to produce a distribution of predicted somatic responses, and compare to the somatic response obtained using the numerical solver with $\text{HoF}_{k}^{\textit{test}}$. Figures 4B and 5B (bottom) show that the voltage traces are very similar, confirming that NOBLE interpolates correctly between models experienced during training, and can be used to obtain new biophysically-realistic models. By contrast, the underlying parameterizations of the HoF models obtained using the multi-objective evolutionary algorithms lack a consistent structure that would allow for meaningful interpolation to discover new biophysically-realistic models. This is exemplified in Figure 5A (bottom) where we see that both voltage responses and F-I curves are completely off when interpolating directly in-between the PDE parameters of the HoF models.
@@ 4.4 Ensemble Predictions
Figure 6: Distributions of voltage traces of $\{\text{HoF}_{i}^{\text{{train}}}\}$ obtained using the numerical solver (left), NOBLE for $\{\text{HoF}_{i}^{\text{{train}}}\}$ (middle), and NOBLE for 200 random samples from the model latent space.
Figure 7: Distributions of key electrophysiological features for ensemble predictions using PDE numerical simulations and synthetic NOBLE predictions when compared to experimental results.
@@
We now exhibit how the single trained NOBLE can be used for ensemble predictions. Given a current injection, we can first run 50 inferences in parallel of NOBLE for the PVALB HoF models $\{\text{HoF}_{i}^{\textit{train}}\}_{i=1}^{50}$ to produce 50 different predicted somatic responses. In Figure 6, we compare these 50 predictions with the corresponding 50 predictions obtained using the numerical solver on $\{\text{HoF}_{i}^{\textit{train}}\}_{i=1}^{50}$. We see that the distribution of curves are very similar, as expected.
@@
Since NOBLE enables interpolation between the HoF models used for training, we can generate novel, biophysically realistic neuron models and are no longer limited to the original 50 HoF models: the trained NOBLE can produce voltage responses for any neuron model whose representation lies within the convex hull $\mathcal{CH}_{\text{{train}}}$ of the training set ${\text{HoF}_{i}^{\text{{train}}}}$. We demonstrate this by querying 200 novel models whose features are sampled randomly within $\mathcal{CH}_{\textit{train}}$. As shown in Figure 6, the distribution of curves remains very similar, but the additional samples provide a denser coverage of the response space while maintaining biophysical realism, with no apparent artifacts or implausible predictions. Figure 7 also shows that the larger distribution of synthetic predictions retains a similar accuracy as the HoF models in terms of the 4 most important electrophysiological features when compared to those of experimental results. These results highlight the capacity of NOBLE to faithfully represent a diverse collection of biophysically realistic neuron models, while allowing for dense interpolation across the model space. As a result, NOBLE could pave the way for improved statistical analysis, and more reliable uncertainty quantification and robust predictive modeling of neuron dynamics.

### v3

@@ 3.2 Current Amplitude Sampling
We consider square DC step current inputs, which are widely used in electrophysiological experiments and common for characterizing neuron behavior. We sample the current amplitudes from a skew-normal distribution whose support matches the experimentally validated range of the HoF models, $I\in[-0.11,0.28]\mathrm{nA}$. To effectively capture the highly nonlinear dynamics around the spiking threshold (0 to $0.05\mathrm{nA}$) where neural responses transition abruptly from being non-spiking to spiking, the mode of our sampling distribution is strategically located within this peri-threshold window. To address the greater learning challenge posed by the high-frequency components of depolarizing, spiking responses (characterized by features such as spike width, latency to first spike, and spike count), we deliberately use a heavier positive tail in our sampling distribution. This ensures the model encounters numerous examples of spike onset and complex spiking patterns during training while still covering the full input range. The distribution of square-pulse amplitudes is shown in Figure˜8.
3.3 Neural Operators for Neuron Dynamics Simulation
We choose to use neural operators as they offer clear advantages for modeling complex dynamics (Section˜C.3). Among neural operators, the Fourier Neural Operator (FNO) is very efficient as it leverages fast Fourier transforms on equidistant grids, which aligns naturally with our setting where both experimental recordings and PDE simulations are sampled at constant timesteps. FNOs provide a principled and efficient framework for modeling neuronal dynamics by learning mappings from input currents to voltage responses across a broad family of neuron models and current injections. Unlike conventional neural networks that operate on vector inputs and outputs of fixed sizes, the FNO learns operators, that is, mappings between functions. By operating in the frequency domain, the FNO efficiently captures global, nonlinear, and high-frequency components of voltage responses. These properties allow the model to generalize across different temporal resolutions, input currents, and neuron types, enabling the accurate simulation of unseen configurations without retraining.
@@ 3.4 Embedding Strategy for Neuron-Model Variability
NOBLE learns a single neural operator that maps from a continuous latent space of user-defined, interpretable neuron characteristics to an ensemble of somatic voltage responses induced by current injection. The frequency-current (F-I) curve is a useful electrophysiological descriptor that summarizes cellular excitability by relating injected current amplitude to the neuron’s firing rate [53]. Differences between HoF parameterizations manifest as shifts in key features of this curve: the threshold current $I_{\textit{thr}}$ (the minimum amplitude that elicits spiking) and the local slope $s_{\textit{thr}}$ at $I_{\textit{thr}}$ (the rate of increase in the firing rate upon spiking). Figure˜3A displays examples of F-I curves for different HoF PVALB models, illustrating how variability in $I_{\textit{thr}}$ and $s_{\textit{thr}}$ can represent the trial-to-trial intrinsic variability observed when a single neuron is repeatedly recorded under the same current injection.
Using this observation, we propose representing a given neuron model by its threshold current $I_{\textit{thr}}$ and local slope $s_{\textit{thr}}$, that is, using $(I_{\textit{thr}},s_{\textit{thr}})$. We propose to use this representation as part of a NeRF-style (Neural Radiance Field) embedding [54], where input features are encoded using sine and cosine functions. More precisely, a feature $p$ is encoded as a stack of trigonometric time-series
$\gamma(p,t)=\left[\sin(2^{0}\pi p\odot t),\ \cos(2^{0}\pi p\odot t),\ \ldots,\ \sin(2^{K-1}\pi p\odot t),\ \cos(2^{K-1}\pi p\odot t)\right],$ (1)
@@
We encode separately the model features $I_{\textit{thr}}$ and $s_{\textit{thr}}$, and the amplitude of the current injection, and stack the resulting embeddings as input channels. To compress the large range of $I_{\textit{thr}}$ and $s_{\textit{thr}}$ values into a manageable scale for embedding, we normalize them to $0.5,3.5]^{2}$, supporting more distinct feature space representations of HoF models. Figure˜10 displays the latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of the 60 HoF PVALB models used in our numerical experiments.
@@ 3.5 NOBLE: Neural Operator with Biologically-informed Latent Embeddings
We introduce the Neural Operator with Biologically-informed Latent Embeddings (NOBLE), for modeling neuronal voltage dynamics in response to current injections. NOBLE offers a scalable alternative to computationally intensive numerical solvers for biophysically detailed, PDE-based neuron models. It learns a direct mapping from input currents and a continuous, interpretable latent space of neuron features, to the resulting voltage traces (Figure˜1A). A key feature of NOBLE is its use of biologically-informed embeddings, which enables interpretability and generalization across biological neuron models. At its core, NOBLE is based on a neural operator, whose discretization invariance allows NOBLE to learn on low-resolution data and infer somatic voltage dynamics at higher resolutions. By combining a neural operator with the proposed continuous interpretable embedding, NOBLE learns a continuous operator over the space of bio-realistic neuron models.
The proposed NOBLE framework offers key advantages that set it apart from existing approaches:
  • NOBLE provides a unified framework that learns ensemble dynamics directly, enabling it to generate diverse, biophysically plausible membrane potentials for the same input. Conditioned on a particular electrophysiological feature, it produces one realization of the intrinsic variability observed in biological neurons. This stands in contrast to previous deep learning approaches, which are inherently deterministic and produce a single trace for each input, failing to capture the trial-to-trial variability observed experimentally. To account for different neuronal behaviors, such models must be retrained for each variation, resulting in inefficiency and fragmentation.
  • Through the latent embedding space of electrophysiological features, NOBLE can interpolate between known HoF models to produce new, bio-realistic neuronal responses. This capability is significant because HoF models are restricted to the finite set discovered by evolutionary optimization, and direct interpolation between their parameters does not result in realistic traces. As shown in Figure˜4, interpolation in NOBLE ’s latent space consistently produces valid, bio-realistic responses, whereas interpolations between PDE parameters do not.
  • NOBLE can rapidly generate arbitrarily many distinct neuron models by sampling points within this continuous latent embedding space and producing the corresponding dynamics. This enables a single model to capture both spiking and subthreshold behaviors beyond the finite set of HoF models, providing an effectively infinite ensemble of bio-realistic responses that remain consistent with the variability observed in biological neurons. This is distinct from previous deep-learning methods, which were limited to predicting either spiking or subthreshold regimes in isolation.
  • The bio-informed latent space of NOBLE, combined with its ability to generate unlimited realistic neuron models, enables fine-grained exploration of neural dynamics. By sampling models across this space, NOBLE can produce somatic responses for different latent features, and reveal how they influence electrophysiological behavior via visualizations (e.g. heat maps and surface plots).
@@ 4.1 Experimental Setup
For evaluating NOBLE, we focus on the PVALB neuron example introduced in Section˜2, and additionally validate generality of the framework on a VIP neuron. Each neuron model $\text{HoF}_{\ell}$ receives as input the stacked embeddings of the normalized model features $(I_{\textit{thr}},s_{\textit{thr}})$ and the applied current injection $I$, and outputs the somatic voltage response. The current injections are square DC steps with an activation duration of $400\mathrm{ms}$, consistent across all stimuli used for training and testing. Given a PVALB model $\text{HoF}_{\ell}$ and a current injection $I$, the neural operator in NOBLE takes the stacked embeddings of the (normalized) model features $I_{\textit{thr}}$ and $s_{\textit{thr}}$ associated to $\text{HoF}_{\ell}$ (with $K=1$ frequency) and $I$ (with $K=9$ different frequencies), and outputs the somatic voltage response. We have access to 60 HoF models, where 50 are used during training, $\{\text{HoF}^{\textit{train}}\}$, and the remaining 10 HoF models $\{\text{HoF}^{\textit{test}}\}$ are used for testing. Figure˜10 displays the latent representations in normalized $(I_{\textit{thr}},s_{\textit{thr}})$-space of these HoF models. For more details on the data generation, see Section˜D.1.
@@
We use a 1D FNO (implemented as in the NeuralOperator library [55]) with $12$ layers, each with $24$ hidden channels and $256$ Fourier modes. The resulting NOBLE with 1.81.8M trainable parameters is trained in PyTorch for $300$ epochs using the Adam optimizer with learning rate $0.004$, and the ReduceLROnPlateau scheduler with factor $0.4$ and patience $4$. The training minimizes the relative L4 error, while the performance metrics are reported using the relative L2 error for interpretability.
To better evaluate physiologically meaningful behavior, we also report errors on four key electrophysiological features: spikecount, AP1_width, mean_AP_amplitude, and steady_state_voltage (see Appendix˜B for definitions of the features). For benchmarking, we compare NOBLE predictions against numerical simulations obtained from HoF models and experimental data since the HoF models were optimized to produce the closest approximations to experimental recordings and capture the biological variability required for a meaningful benchmark. Further details on the evaluation and evaluation metrics are provided in Section˜D.2.
The PyTorch codes used for our implementation of NOBLE and the numerical experiments are based on the NeuralOperator library [55], and are made available at github.com/neuraloperator/noble.
4.2 Testing on HoF Models Included in the Training Set
We first validate that the trained NOBLE can accurately reproduce the somatic voltage responses of the training $\{\text{HoF}^{\textit{train}}\}$ models when tested on current injections not seen during training. Figure˜3C shows that the voltage traces exhibit minimal differences, confirming that NOBLE generalizes well to unseen inputs. This is supported by a relative L2 test error of $2.18\%$ with the $\{\text{HoF}^{\textit{train}}\}$ models. Figure˜3B also shows that the numerical solver outputs align closely with experimental recordings, and together with Figure˜3C, indicates that NOBLE inherits this agreement and captures physiologically meaningful dynamics. Note that the available experimental recordings correspond to stimuli with activation durations of 1s. NOBLE also achieves errors of $3\%$ for spikecount, $32.8\%$ for AP1_width, $8.54\%$ for mean_AP_amplitude, and $0.83\%$ for steady_state_voltage. To further assess NOBLE’s applicability across neuron types, we trained it on a VIP neuron using the same architecture and observed similarly strong performance (see Section˜D.3 for more details).
The relatively higher error on AP1_width arises from how the feature is computed: it measures the width of the first spike at half amplitude, where the half level is defined between the spike peak and the subsequent after-hyperpolarization minimum. If the predicted peak or minimum is slightly misaligned relative to the ground truth, the half-voltage reference shifts, and the measured width corresponds to a different portion of the trace. This sensitivity makes AP1_width less stable to small deviations, so its relative error should be interpreted with caution compared to the other features.
@@ 4.2 Testing on HoF Models Included in the Training Set
We also generate the F-I curves using the trained NOBLE for $\{\text{HoF}^{\textit{train}}\}$ and compare them with the reference F-I curves produced by the numerical solver for the same HoF models. As shown in Figure˜3A, the curves from both methods remain close overall, although for 3 out of the 50 HoF models the NOBLE predictions show larger deviations in firing rate between $0.0-0.1\mathrm{nA}$.
Figure 3: A) F–I curves from experimental recordings, PDE simulations, and NOBLE predictions on $\{\text{HoF}^{\textit{train}}\}$. For one HoF model, B) compares experimental voltage responses with PDE simulations at a current injection of $0.1\mathrm{nA}$ (top) and $0.11\mathrm{nA}$ (bottom), and C) compares the corresponding PDE simulation with the NOBLE prediction for the same HoF model.
@@ 4.3 Interpolating Between Models
Figure 4: A) F–I curves from experimental recordings, PDE simulations, and NOBLE predictions on 50 interpolated HoF models. For one interpolated model, B) compares experimental voltage responses with PDE simulations at a current injection of $0.1\mathrm{nA}$, and C) compares the corresponding PDE simulation with the NOBLE prediction for the same interpolated HoF model.
We now test NOBLE’s ability to interpolate between HoF models, within the convex hull $\mathcal{CH}_{\textit{train}}$ of the $\{\text{HoF}^{\textit{train}}\}$ models used for training. Given a model $\text{HoF}_{k}^{\textit{test}}$ excluded during training, we randomly sample 50 points in the model latent space close to the defining $(I_{\textit{thr}},s_{\textit{thr}})$ features of $\text{HoF}_{k}^{\textit{test}}$ (Figure˜11 illustrates the small circle around $\text{HoF}_{k}^{\textit{test}}$ in which the 50 synthetic models are sampled), none of which has been experienced during training. We run NOBLE in parallel on these sampled models to generate the corresponding somatic voltage responses and compare them to the response obtained from the numerical solver with $\text{HoF}_{k}^{\textit{test}}$. Figure˜4C shows that, when interpolating between models within the latent space, NOBLE accurately captures the underlying neuronal dynamics and produces voltage responses consistent with experimental recordings. In contrast, Figure˜4B shows that the underlying parameterizations of the HoF models obtained using the multi-objective evolutionary algorithms lack a consistent structure that would allow for meaningful interpolation to discover new bio-realistic models. This difference is further illustrated in Figure˜4A, where the F-I curves generated by the PDE models deviate markedly from experimental data, while those produced by NOBLE remain biophysically meaningful and closely aligned with observed neuronal behavior.
[…]
@@ 4.4 Ensemble Predictions
We now examine how the single trained NOBLE can be used for ensemble predictions. Given a current injection, we run 50 inferences in parallel of NOBLE for the $\{\text{HoF}^{\textit{train}}\}$ models to produce 50 somatic responses. In Figure˜5A (left, middle), we compare these 50 predictions with the corresponding 50 numerical solver simulations from $\{\text{HoF}^{\textit{train}}\}$. We see that the distribution of curves is very similar.
@@
Since NOBLE enables interpolation between the HoF models used for training, it can generate novel bio-realistic neuron models and produce voltage responses for any neuron model whose latent space representation lies within the convex hull $\mathcal{CH}_{\textit{train}}$ of the training set $\{\text{HoF}^{\textit{train}}\}$. We demonstrate this by querying 200 novel models whose features are sampled randomly within $\mathcal{CH}_{\textit{train}}$. Figure˜5A (right) shows that the distribution of curves remains very similar, but the additional samples provide a denser coverage of the response space while maintaining bio-realism, with no artifacts or implausible predictions. We include additional ensemble analyses in Section˜D.5. There, we first validate NOBLE on the test models $\{\text{HoF}^{\textit{test}}\}$, where predicted and ground-truth responses again show close agreement (Figure˜12). Second, we examine local perturbations in the latent embedding space by sampling 50 synthetic models from a small circle around a held-out test model (Figure˜11). The resulting responses (Figure˜13) resemble small perturbations around the unseen ground-truth trace, demonstrating NOBLE’s ability to generalize to unseen models and the smoothness of the latent space.
Figure 5: A) Distributions of somatic voltage traces across HoF models for current injections of $0.1\mathrm{nA}$ (top) and $0.11\mathrm{nA}$ (bottom). B) Relative errors of ensemble predictions from PDE simulations and NOBLE models on $\{\text{HoF}^{\textit{train}}\}$ compared to experimental recordings across the four key features.
Figure˜5B shows that NOBLE’s ensemble predictions for the full collection of HoF models achieve comparable accuracy to the HoF models themselves when evaluated on the four key electrophysiological features relative to experimental data. Here, we report the mean_frequency, representing the average firing rate, to account for the difference in stimulus activation durations between the experimental recordings and NOBLE, since spikecount is a nonlinear function of time and cannot be directly rescaled. These results demonstrate NOBLE’s ability to faithfully represent a diverse set of bio-realistic neuron models while enabling dense interpolation across the model space. However, experimental recordings for this neuron are limited to a single trace across nine amplitudes. Consequently, our comparison is against a single realization rather than a distribution. In practice, experimental features exhibit trial-to-trial variability, which NOBLE is designed to capture, but cannot be directly validated here due to limited experimental data availability.
4.5 Choice of Biologically-Informed Latent Embeddings
Feature embeddings are central to NOBLE as they enable generalization across and in-between biological neuron models. Here, we have chosen $I_{thr}$ and $s_{thr}$ for their strong biological interpretability, as a natural 2D representation of firing and non-firing dynamics. To quantify the contribution of embeddings, we conduct an ablation study evaluating NOBLE with lower-dimensional embeddings, as well as without embeddings. We also consider a higher-dimensional embedding that includes AHP_depth, selected for its large variation across intracell HoF models and its low correlation with $I_{thr}$ and $s_{thr}$. Results are reported in Table˜1 in terms of the relative L2 error of predicted voltage traces and the four key electrophysiological features on the test set with $\{\text{HoF}^{\textit{train}}\}$ models.
Without embeddings, NOBLE fails to predict any firing responses, indicating that embeddings are necessary to capture spiking behavior. With a single embedded feature, either $s_{thr}$ or $I_{thr}$, NOBLE achieves similar accuracy on voltage traces and steady_state_voltage, but errors on spike-related features remain large, with $I_{thr}$ providing better accuracy for AP1_width. As discussed earlier, this feature is particularly sensitive to small misalignments in spike peak and after-hyperpolarization minima, which makes its relative error less robust as a metric. Embedding both $s_{thr}$ and $I_{thr}$ yields the best overall performance across all features as well as the voltage trace. Extending the embedding space with AHP_depth slightly improves AP1_width, but reduces accuracy for the other features. These results show that feature embeddings are important for NOBLE to capture both firing and non-firing dynamics, and more broadly for its ability to generalize across diverse neuron models.
Table 1: Relative L2 error of NOBLE on voltage traces and the four key features, when trained with different embedded features. Results are evaluated on the test set with $\{\text{HoF}^{\textit{train}}\}$ models.
Features embedded Voltage Steady state voltage Spikecount AP1 width Mean AP amplitude
None 12.1% 1.31% Never fires Never fires Never fires
$s_{thr}$ 2.83% 1.33% 4.9% 233% 13%
$I_{thr}$ 2.73% 1.20% 4.4% 107% 14%
$s_{thr},I_{thr}$ 1.92% 1.02% 3.1% 27% 8.9%
$s_{thr},I_{thr},$ AHP_depth 2.16% 1.04% 3.3% 22% 9.5%
While we only considered constructing the latent space from electrophysiological features, NOBLE can readily incorporate additional modalities such as gene expression profiles from patch-sequencing data [56]. Learning joint embeddings across modalities could yield a unified latent space linking gene expression, electrophysiology, and morphology, providing a means to test hypotheses that are infeasible experimentally, such as how genes associated with neurological diseases [49] influence neuronal dynamics. In doing so, NOBLE would pave the way for improved statistical analysis, more reliable uncertainty quantification, and robust predictive modeling of neuronal behavior.
4.6 Feature Specific Physics-Informed Fine-Tuning of NOBLE
To preserve overall neural dynamics while improving accuracy on a single specific feature, NOBLE can be fine-tuned with a weighted composite loss $\mathcal{L}(\lambda)=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{F}$, where $\mathcal{L}_{F}$ penalizes deviations in feature $F$, and $\lambda$ controls its influence. We illustrate this by fine-tuning NOBLE to improve sag_amplitude accuracy, a feature reflecting the hyperpolarization-activated cation channel (Ih). In the human cortex, Ih expression varies with cortical depth, making sag_amplitude a relevant physiological marker. We fine-tune on 19,600 subthreshold stimuli with negative amplitudes. Even without a feature-specific loss, fine-tuning reduces the L2 feature error from 70% to 19.2%, and incorporating $\mathcal{L}_{F}$ further lowers it to 9.6% while preserving overall signal fidelity. These results show that NOBLE can be selectively refined to prioritize biophysical features of interest without compromising overall performance. Further details are provided in Section˜D.7.